# Coherence-Mediated Quantum Thermometry in a Hybrid Circuit-QED Architecture

### v1

@@ V Outlook and Conclusion
The principle demonstrated here, using a coherent ancillary mode as an information buffer between a noisy environment and a fragile qubit, defines a general paradigm for quantum-enhanced thermometry and noise spectroscopy. In practice, the scheme is compatible with state-of-the-art circuit-QED hardware, where high-$Q$ 3D cavities routinely provide interaction times of tens of milliseconds and cross-Kerr couplings in the 10105050 kHz range can be engineered with multi-qubit pathways. Beyond temperature estimation, related strategies may be adapted to quantum calorimetry, fluctuation spectroscopy, or the characterization of correlated noise in quantum processors. More broadly, coherence-mediated sensing architectures could provide useful tools in nuclear and high-energy physics experiments, where ultrasensitive calorimetry at millikelvin scales is urgently needed. Taken together, these considerations point to broad opportunities for coherence-mediated sensing, while also underscoring the immediate impact of our results for quantum thermometry.
In conclusion, we have introduced and analyzed a coherence-mediated thermometry scheme in a hybrid cQED platform, where a thermal mode is monitored indirectly through its cross-Kerr coupling to a coherent probe and read out by a dispersively coupled qubit. By separating information acquisition from readout, the scheme circumvents the intrinsic dephasing bottleneck of qubit-only thermometers, enabling longer interaction times, higher sensitivity, and greater robustness. These results establish qubit-mediated cross-Kerr thermometry as a fundamentally distinct and scalable approach, providing a path toward nondissipative, reusable quantum thermometers with broad applications across quantum science and technology.
@@ Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract no. DE-AC02-07CH11359.
@@ Appendix A QFI from a Coherence Envelope
$\rho(T)=\frac{1}{2}\begin{pmatrix}1&C(T)\,e^{i\Phi(T)}\\ C(T)\,e^{-i\Phi(T)}&1\end{pmatrix}$, (11)
where $C(T)$ denotes the coherence envelope, i.e. the visibility of Ramsey fringes, and $\Phi(T)$ is a mean accumulated phase. Both quantities generally depend on the interaction time $\tau$; here we suppress this dependence for brevity.
Physically, $C(T)$ captures the loss of coherence induced by temperature-dependent fluctuations (e.g. photon-number noise in a coupled mode), while $\Phi(T)$ represents a deterministic phase shift that can also carry temperature information. In this sense, $C(T)$ quantifies the amplitude of the Bloch vector and $\Phi(T)$ its orientation in the equatorial plane of the Bloch sphere.
@@
$\mathbf{r}(T)=\big(C\cos\Phi,\,C\sin\Phi,\,0\big),\qquad r=|\mathbf{r}|=C.$
For a general qubit state $\rho=\tfrac{1}{2}(\mathbb{I}+\mathbf{r}\cdot\bm{\sigma})$, the quantum Fisher information (QFI) for estimating a parameter $T$ is known to be [19, 20]
$\mathcal{F}_{T}=\frac{\big(\partial_{T}r\big)^{2}}{1-r^{2}}+r^{2}\big(\partial_{T}\phi\big)^{2},\quad\text{with}\;r=C,\;\phi=\Phi.$ (12)
This expression separates the two distinct resources available for thermometry: changes in the coherence amplitude $C(T)$ and changes in the phase angle $\Phi(T)$.
Explicitly,
@@
Although the above derivation was formulated for a qubit undergoing temperature-dependent dephasing, the resulting expression is fully general for any coherence channel characterized by a complex visibility $C(T)e^{i\Phi(T)}$. In the main text, this formalism is applied to the probe mode: the absorber’s temperature modifies the probe’s complex amplitude in the same way, with $C(T)$ describing the loss of visibility and $\Phi(T)$ the temperature-dependent phase shift. The qubit in that scheme serves only as a phase-sensitive transducer that converts the probe’s coherence into a measurable Ramsey signal. Therefore, Eq. (13) quantifies the fundamental temperature information accessible in both the qubit- and probe-based representations.
@@ Appendix B Derivation of the Coherence-Mediated Envelope and QFI
$|\alpha\rangle\;\longrightarrow\;|\alpha e^{i\phi_{b}}\rangle,\qquad\phi_{b}=\lambda n_{a}\tau.$
@@
For a thermal state, $n_{a}$ follows a Bose–Einstein distribution with mean $\bar{n}_{a}(T)$ and variance $\mathrm{Var}(n_{a})=\bar{n}_{a}\big(1+\bar{n}_{a}\big)$. The resulting phase variance is
$\sigma_{\phi}^{2}=\mathrm{Var}(\phi)=(\lambda\tau)^{2}\bar{n}_{a}(T)\big[1+\bar{n}_{a}(T)\big].$ (16)
@@
The qubit coherence envelope is proportional to the overlap of the original probe state with the diffused mixture:
@@
Using the coherent-state overlap $\langle\alpha|\beta\rangle=\exp\!\left[\,\alpha^{*}\beta-\tfrac{1}{2}\big(|\alpha|^{2}+|\beta|^{2}\big)\right],$ with $\beta=\alpha e^{i\phi}$, we obtain
$|\langle\alpha|\alpha e^{i\phi}\rangle|^{2}=\exp\!\big[-2|\alpha|^{2}(1-\cos\phi)\big]=e^{-2\alpha^{2}}e^{2\alpha^{2}\cos\phi}.$
@@
$C(\tau,T)=\exp\!\left[-2\alpha^{2}\left(1-e^{-\sigma_{\phi}^{2}}\right)\right].$ (20)
This expression identifies the effective dephasing strength
$\Gamma_{\phi}(T)=\sigma_{\phi}^{2}=(\lambda\tau)^{2}\bar{n}_{a}(T)\big[1+\bar{n}_{a}(T)\big],$ (21)
which reduces to $\Gamma_{\phi}(T)\approx(\lambda\tau)^{2}\bar{n}_{a}(T)$ for $\bar{n}_{a}\ll 1$. The final coherence envelope is therefore
$C(\tau,T)=\exp\!\left[-2\alpha^{2}\Big(1-e^{-\Gamma_{\phi}(T)}\Big)\right].$ (22)
At low temperatures $\bar{n}_{a}(T)\to 0$], the dephasing vanishes and $C(\tau,T)\to 1$. At high temperatures, the envelope saturates to $\exp(-2\alpha^{2})$, corresponding to complete phase scrambling. The nonlinear dependence on both $\alpha$ and $\bar{n}_{a}$ makes the scheme sensitive to thermal fluctuations even deep in the sub-photon regime. Importantly, unlike the qubit-only protocol where dephasing occurs directly on the qubit, here the qubit can remain dynamically decoupled while the probe collects thermal information. This distinction underlies the extended interaction times achievable in the coherence-mediated scheme when the probe is implemented as a high-$Q$ cavity.
In the coherence-mediated protocol, all $T$-dependence arises from the envelope $C(\tau,T)$; there is no temperature-dependent phase offset $\Phi(T)$. Differentiating Eq. (22) gives
$\partial_{T}C=C\cdot 2\alpha^{2}e^{-\Gamma_{\phi}(T)}\,\Gamma_{\phi}^{\prime}(T),$ (23)
with
$\Gamma_{\phi}^{\prime}(T)=(\lambda\tau)^{2}\partial_{T}\bar{n}_{a},\qquad\partial_{T}\bar{n}_{a}=\frac{\hbar\omega_{a}}{k_{B}T^{2}}\,\bar{n}_{a}(T)\big[1+\bar{n}_{a}(T)\big].$
@@
$\mathcal{F}_{T}(\tau,T)=\frac{\Big[2\alpha^{2}e^{-\Gamma_{\phi}(T)}(\lambda\tau)^{2}\,\partial_{T}\bar{n}_{a}\Big]^{2}\,C(\tau,T)^{2}}{1-C(\tau,T)^{2}}.$ (24)
In the low-occupation limit $\bar{n}_{a}\ll 1$ (so $\Gamma_{\phi}\ll 1$ and $C\simeq e^{-2\alpha^{2}(\lambda\tau)^{2}\bar{n}_{a}}$), this simplifies to
$\mathcal{F}_{T}(\tau,T)\;\approx\;\alpha^{2}(\lambda\tau)^{2}\left(\frac{\hbar\omega_{a}}{k_{B}T^{2}}\right)^{2}\bar{n}_{a}(T).$ (25)
This shows that the QFI scales linearly with thermal occupation $\bar{n}_{a}(T)$, quadratically with the effective interaction strength $\lambda\tau$, and linearly with the probe photon number $|\alpha|^{2}$, highlighting the bosonic amplification inherent in the coherence-mediated approach.
Appendix C Probe Phase-Shift QFI
@@
The temperature-dependent phase shift transforms an initial coherent probe $|\alpha\rangle$ into
$|\psi(T)\rangle=|\alpha e^{i\phi_{b}(T)}\rangle.$ (27)
The quantum Fisher information (QFI) for temperature estimation from this pure state is [17, 53, 54]
@@
$\mathcal{F}_{\Phi}(T)=4|\alpha|^{2}\big(\partial_{T}\phi_{b}(T)\big)^{2}.$ (29)
Thus, the thermometric sensitivity of the heterodyne phase-tracking scheme grows quadratically with the temperature derivative of the induced phase. Although increasing intracavity photons $|\alpha|^{2}$ enhances $\mathcal{F}_{\Phi}$ linearly, it introduces potential backaction and power-induced dephasing. To remain in the weak-measurement regime, we set $|\alpha|^{2}=1$ in the following expressions. We also choose to absorb the factor of $4$ into the operational sensitivity definition used for comparison across schemes.
Finally, inserting the explicit dependence of $\phi_{b}(T)=\lambda\tau\bar{n}_{a}$ gives the compact form
$\mathcal{F}_{\Phi}(T)=\big(\lambda\tau\,\partial_{T}\bar{n}_{a}\big)^{2},$ (30)
which quantifies the thermometric sensitivity of the phase-shift scheme in the single-photon normalization.
@@ Appendix D Derivation of Qubit-Only-Based Coherence Envelope and QFI
[…]
We consider a qubit dispersively coupled to a single harmonic mode $\hat{a}$ of frequency $\omega_{a}$ and linewidth $\kappa_{a}$. In the dispersive regime the interaction takes the longitudinal form
$\hat{H}_{\mathrm{int}}=\chi_{a}\,\hat{a}^{\dagger}\hat{a}\,\hat{\sigma}_{z},$
so that the qubit transition frequency is shifted by $\pm\chi_{a}n$ conditional on the photon number $n$ (cross-Kerr shift $\chi_{a}$).
Photon-number fluctuations of the thermal mode are modeled as stationary Gaussian noise with exponential correlations,
$C_{nn}(\tau)\equiv\big\langle\delta\hat{n}(\tau)\,\delta\hat{n}(0)\big\rangle=\bar{n}(\bar{n}+1)\,e^{-\kappa_{a}|\tau|},\qquad\delta\hat{n}\equiv\hat{a}^{\dagger}\hat{a}-\bar{n},$
where $\bar{n}=[\exp(\hbar\omega_{a}/k_{B}T)-1]^{-1}$. Equivalently, the symmetrized noise spectrum is
$S_{nn}(\omega)=\int_{-\infty}^{\infty}d\tau\,e^{i\omega\tau}\,C_{nn}(\tau)=\frac{2\kappa_{a}\,\bar{n}(\bar{n}+1)}{\kappa_{a}^{2}+\omega^{2}}.$ (31)
During a Ramsey sequence of duration $\tau$, the qubit accumulates a stochastic phase
$\phi(\tau)=2\chi_{a}\int_{0}^{\tau}\!dt\,\hat{n}(t)=2\chi_{a}\bar{n}\,\tau+2\chi_{a}\int_{0}^{\tau}\!dt\,\delta\hat{n}(t).$ (32)
Averaging over Gaussian fluctuations using the cumulant expansion, the off-diagonal coherence is
$C_{q}(\tau)=\frac{\rho_{eg}(\tau)}{\rho_{eg}(0)}=e^{-i2\chi_{a}\bar{n}\tau}\;\exp\!\left[-\tfrac{1}{2}\big\langle\!\big\langle\phi_{\mathrm{noise}}^{2}(\tau)\big\rangle\!\big\rangle\right],$
@@
$\big\langle\!\big\langle\phi_{\mathrm{noise}}^{2}(\tau)\big\rangle\!\big\rangle=(2\chi_{a})^{2}\int_{0}^{\tau}\!\!\int_{0}^{\tau}dt_{1}\,dt_{2}\,C_{nn}(t_{1}-t_{2}).$
@@
$\int_{0}^{\tau}\!\!\int_{0}^{\tau}\!dt_{1}\,dt_{2}\,e^{-\kappa_{a}|t_{1}-t_{2}|}=2\int_{0}^{\tau}\!d\theta\,(\tau-\theta)\,e^{-\kappa_{a}\theta}=\frac{2\big(\kappa_{a}\tau-1+e^{-\kappa_{a}\tau}\big)}{\kappa_{a}^{2}}.$

### v2

@@ V Outlook and Conclusion
The principle demonstrated here, using a coherent mode as an information buffer between a noisy environment and a fragile qubit, provides a key advantage by protecting the qubit’s coherence while preserving temperature information in the buffered mode, thereby defining a general paradigm for coherence-mediated thermometry and noise spectroscopy. The regime of advantage for CMQT is the ultra-low-occupation limit $\bar{n}_{a}\!\ll\!1$, where conventional dispersive thermometry is fundamentally signal-limited; at higher temperatures the marginal benefit diminishes and simpler approaches can suffice.
In practice, the approach is fully compatible with state-of-the-art cQED hardware. The high-$Q$ 3D cavity can support millisecond coherence times, and the moderate cross-Kerr coupling can be engineered via multi-qubit pathways. Beyond temperature estimation, closely related strategies can be extended to quantum calorimetry, fluctuation spectroscopy, and the characterization of correlated noise in superconducting quantum processors. A natural next step is to cast the protocol in a multiparameter estimation framework, treating $T$ together with nuisance parameters such as effective absorber coupling rates, residual dispersive pulls, or detection efficiency; this would enable explicit robustness bounds and optimal experimental design under realistic calibration uncertainty. More broadly, coherence-mediated sensing architectures offer a promising route for ultrasensitive calorimetry in nuclear and high-energy physics experiments, where millikelvin-scale energy resolution is increasingly critical. Taken together, these considerations highlight both the broad applicability of coherence-mediated sensing and the immediate impact of our results for quantum thermometry.
In conclusion, we have introduced and analyzed a coherence-mediated thermometry scheme within a hybrid cQED architecture, in which thermal fluctuations are monitored indirectly through its cross-Kerr interaction with a coherent probe and read out via a dispersively coupled qubit. By cleanly separating information acquisition from readout, the scheme circumvents the intrinsic self-dephasing bottleneck of qubit-based thermometers, enabling longer interaction times, higher per-shot sensitivity, and substantially improved robustness. These results establish coherence-mediated quantum thermometry as a fundamentally distinct and scalable strategy, opening a path toward nondissipative, reusable quantum metrology with broad applicability across quantum science, precision sensing, and low-temperature detector technologies.
The data that support the findings of this article are openly available [15].
@@ Acknowledgements.
This work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS), under Contract No. 89243024CSC000002. Fermilab is operated by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.
@@ Appendix A QFI from a Coherence Envelope
$\rho(T)=\frac{1}{2}\begin{pmatrix}1&C\,e^{i\Phi}\\ C\,e^{-i\Phi}&1\end{pmatrix}$, (10)
where the temperature-dependent $C$ and $\Phi$ denote, respectively, the coherence envelope (the visibility of Ramsey fringes) and the mean accumulated phase. Both quantities generally depend on the interaction time $\tau$; here we suppress this dependence for brevity.
Physically, $C$ captures the loss of coherence induced by temperature-dependent fluctuations (e.g. photon-number noise in a coupled mode), while $\Phi$ represents a deterministic phase shift that can also carry temperature information. In this sense, $C$ quantifies the amplitude of the Bloch vector and $\Phi$ its orientation in the equatorial plane of the Bloch sphere.
@@
$\mathbf{r}(T)=\big(C\cos\Phi,\,C\sin\Phi,\,0\big),\quad r=|\mathbf{r}|=C.$
For a general qubit state $\rho=\tfrac{1}{2}(\mathbb{I}+\mathbf{r}\cdot\boldsymbol{\sigma})$, the quantum Fisher information (QFI) for estimating a parameter $T$ is known to be [40, 67]
$\mathcal{F}_{T}=\frac{\big(\partial_{T}r\big)^{2}}{1-r^{2}}+r^{2}\big(\partial_{T}\phi\big)^{2},$ (11)
with $r=C,\mathrm{and}\phi=\Phi$. This expression separates the two distinct resources available for thermometry: changes in the coherence amplitude $C$ and changes in the phase angle $\Phi$. These are two complementary information channels in the likelihood: amplitude/visibility changes and phase shifts. Explicitly,
@@
Although the above derivation was formulated for a qubit undergoing temperature-dependent dephasing, the resulting expression is fully general for any coherence channel characterized by a complex visibility $Ce^{i\Phi}$. In the main text, this formalism is applied to the probe mode: the absorber’s temperature modifies the probe’s complex amplitude in the same way, with $C$ describing the loss of visibility and $\Phi$ the temperature-dependent phase shift. The qubit in that scheme serves only as a phase-sensitive transducer that converts the probe’s coherence into a measurable Ramsey signal. Therefore, Eq. (12) quantifies the fundamental temperature information accessible in both the qubit- and probe-based representations.
@@ Appendix B Derivation of the Coherence-Mediated Envelope and QFI
$|\alpha\rangle\;\rightarrow\;|\alpha e^{i\phi_{b}}\rangle,\quad\phi_{b}=\lambda n_{a}\tau.$
@@
For a thermal state, $n_{a}$ follows a Bose–Einstein distribution with mean $\bar{n}_{a}$ and variance $\mathrm{Var}(n_{a})=\bar{n}_{a}(1+\bar{n}_{a})$. The resulting phase variance is
$\sigma_{\phi}^{2}=\mathrm{Var}(\phi)=(\lambda\tau)^{2}\bar{n}_{a}(1+\bar{n}_{a}).$ (15)
@@
The probe coherence envelope is proportional to the overlap of the original probe state with the diffused mixture:
@@
We obtain
$|\langle\alpha|\alpha e^{i\phi}\rangle|^{2}$ $=\exp\!\big[-2|\alpha|^{2}(1-\cos\phi)\big]$
$=e^{-2\alpha^{2}}e^{2\alpha^{2}\cos\phi}.$
using $\langle\alpha|\beta\rangle=\exp\!\left[\,\alpha^{*}\beta-\tfrac{1}{2}\big(|\alpha|^{2}+|\beta|^{2}\big)\right]$ and $\beta=\alpha e^{i\phi}$.
@@
$C(\tau,T)=\exp\!\left[-2\alpha^{2}\left(1-e^{-\sigma_{\phi}^{2}/2}\right)\right].$ (19)
We define the effective dephasing strength as
$\Gamma_{\phi}(T)=\sigma_{\phi}^{2}/2=\frac{1}{2}(\lambda\tau)^{2}\bar{n}_{a}(1+\bar{n}_{a}).$ (20)
The final coherence envelope is therefore
$C(\tau,T)=\exp\!\big[-2\alpha^{2}(1-e^{-\Gamma_{\phi}})\big].$ (21)
At low temperatures ($\bar{n}_{a}\to 0$), the dephasing vanishes and $C(\tau,T)\to 1$. At high temperatures, the envelope saturates to $\exp(-2\alpha^{2})$, corresponding to complete phase scrambling. The nonlinear dependence on both $\alpha$ and $\bar{n}_{a}$ makes the scheme sensitive to thermal fluctuations even deep in the sub-photon regime.
In the CM protocol, all $T$-dependence arises from the envelope $C(\tau,T)$; there is no temperature-dependent phase offset $\Phi(T)$. Differentiating Eq. (21) gives
$\partial_{T}C=C\cdot 2\alpha^{2}e^{-\Gamma_{\phi}}\,\partial_{T}\Gamma_{\phi},$ (22)
with $\partial_{T}\Gamma_{\phi}=\tfrac{1}{2}(\lambda\tau)^{2}(1+2\bar{n}_{a})\partial_{T}\bar{n}_{a},$ and $\partial_{T}\bar{n}_{a}=\hbar\omega_{a}\,\bar{n}_{a}(1+\bar{n}_{a})/k_{B}T^{2}.$
@@
$\mathcal{F}_{T}(\tau,T)=\frac{\Big[\alpha^{2}e^{-\Gamma_{\phi}}(\lambda\tau)^{2}\,(1+2\bar{n}_{a})\,\partial_{T}\bar{n}_{a}\Big]^{2}\,C^{2}}{1-C^{2}}.$ (23)
In the low-occupation and weak-dephasing limit ($\bar{n}_{a}\ll 1$ and $\Gamma_{\phi}\ll 1$), Eq. (23) simplifies to
$\mathcal{F}_{T}(\tau,T)\;\approx\;\alpha^{2}(\lambda\tau)^{2}\left(\frac{\hbar\omega_{a}}{k_{B}T^{2}}\right)^{2}\bar{n}_{a}.$ (24)
This shows that the QFI scales linearly with thermal occupation $\bar{n}_{a}(T)$, quadratically with the effective interaction strength $\lambda\tau$, and linearly with the probe photon number $|\alpha|^{2}$, highlighting the bosonic amplification inherent in the CM approach.
Appendix C Derivation of Phase-Shift QFI
@@
The temperature-dependent phase shift transforms an initial coherent probe $|\alpha\rangle$ into $|\psi(T)\rangle=|\alpha e^{i\phi_{b}(T)}\rangle.$ The quantum Fisher information (QFI) for temperature estimation from this pure state is [46, 5, 61]
@@
$\mathcal{F}_{\Phi}(T)=4|\alpha|^{2}\big(\partial_{T}\phi_{b}(T)\big)^{2}=4|\alpha|^{2}\big(\lambda\tau\,\partial_{T}\bar{n}_{a}\big)^{2}.$ (27)
Equation (27) shows that the PS strategy benefits from a simple, fully coherent mapping: the temperature dependence enters only through the mean phase $\phi_{b}(T)$, yielding a clean $|\alpha|^{2}\tau^{2}$ scaling that can be boosted by increasing probe power or interaction time. At the same time, the sensitivity diminishes at low temperature through the factor $\partial_{T}\bar{n}_{a}$, reflecting the reduced thermal susceptibility of the absorber mode. Thus, this scheme offers a straightforward and hardware–efficient thermometric readout, albeit with a narrower useful window than coherence-based sensing.
@@ Appendix D Derivation of Qubit-Only-Based Coherence Envelope and QFI
Figure 7: (a) QFI $\mathcal{F}_{q}(T,\tau)$ as a function of temperature $T$ and interaction time $\tau$ for a representative dispersive coupling $\chi_{a}/2\pi=50~\mathrm{kHz}$. The maximum $\mathcal{F}_{q}$ and corresponding optimal $\tau^{\star}$ both decrease as the thermal occupation increases (white dashed line). (b) $\mathcal{F}_{q}$ versus $\tau$ for fixed absorber linewidths $\kappa_{a}/2\pi=10,\,10^{2},\,10^{3}~\mathrm{kHz}$ at $T=10~\mathrm{mK}$. Small $\kappa_{a}$ (quasi-static limit) yields a finite optimal $\tau^{\star}$, while large $\kappa_{a}$ produces Markovian dynamics for which $\mathcal{F}_{q}(\tau)$ grows monotonically. (c) $\mathcal{F}_{q}$ versus $\tau$ for different dispersive couplings $\chi_{a}/2\pi=10,\,100~\mathrm{kHz}$. Increasing $\chi_{a}$ compresses the optimal sensing window to shorter $\tau^{\star}$, while the peak value of $\mathcal{F}_{q}$ remains nearly unchanged, indicating that stronger coupling mainly reshapes the per-shot information rather than enhances it.
We consider a qubit dispersively coupled to a single harmonic mode $\hat{a}$ with frequency $\omega_{a}$ and linewidth $\kappa_{a}$. To maintain consistency with the definition used in the CM scheme, we write the dispersive interaction in longitudinal form
$\hat{H}_{\mathrm{int}}=\tfrac{1}{2}\,\chi_{a}\,\hat{a}^{\dagger}\hat{a}\,\hat{\sigma}_{z},$
so that $\chi_{a}$ denotes the qubit frequency shift per thermal photon.
For thermometry, the relevant effect of the thermal mode is the stochastic modulation of its photon number. We treat $\delta\hat{n}_{a}(t)=\hat{a}^{\dagger}(t)\hat{a}(t)-\bar{n}_{a}$ as a stationary random process fully characterized by its two-point correlator. In the weak-dispersive and weak-dephasing regime, the qubit’s phase evolution is accurately captured by the second-order cumulant (Gaussian-phase) approximation [11, 25, 4]. We define the thermal correlation as
$C_{nn}(\tau)=\langle\delta\hat{n}_{a}(\tau)\,\delta\hat{n}_{a}(0)\rangle=\bar{n}_{a}(\bar{n}_{a}+1)\,e^{-\kappa_{a}|\tau|}.$ (28)
During a Ramsey evolution of duration $\tau$, the qubit accumulates the stochastic phase
$\phi(\tau)=\chi_{a}\!\int_{0}^{\tau}dt\,\hat{n}_{a}(t)=\chi_{a}\bar{n}_{a}\,\tau+\chi_{a}\!\int_{0}^{\tau}dt\,\delta\hat{n}_{a}(t).$ (29)
The cumulant expansion yields the coherence
$C_{q}(\tau)=\exp[-i\chi_{a}\bar{n}_{a}\,\tau]\,\exp\!\Big[-\tfrac{1}{2}\,\langle\!\langle\phi_{\rm noise}^{2}(\tau)\rangle\!\rangle\Big],$ (30)
@@
$\langle\!\langle\phi_{\rm noise}^{2}(\tau)\rangle\!\rangle=\chi_{a}^{2}\int_{0}^{\tau}\!\!\int_{0}^{\tau}dt_{1}\,dt_{2}\,C_{nn}(t_{1}-t_{2}).$
@@
$\int_{0}^{\tau}\!\!\int_{0}^{\tau}dt_{1}dt_{2}\,e^{-\kappa_{a}|t_{1}-t_{2}|}=\frac{2(\kappa_{a}\tau-1+e^{-\kappa_{a}\tau})}{\kappa_{a}^{2}}.$